OpenStudy (anonymous):

what is k-5/12=negative 4/15

OpenStudy (anonymous):

the 4 is negative

OpenStudy (anonymous):

So its k-5/12=-4/15

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

is it -4/15 or -4/12 ?

OpenStudy (phi):

can you simplify the left side ?

OpenStudy (anonymous):

yes into 10/15

OpenStudy (phi):

-5/12 + 5/12 is ?

OpenStudy (anonymous):

no the 4 is the only negative number

OpenStudy (phi):

oh. can you use the equation editor at the bottom to put in the problem? I am lost.

OpenStudy (anonymous):

ok \[k-5/12= -4/12\]

OpenStudy (anonymous):

whats the answer

OpenStudy (phi):

Is it -4/12 on the right side? It was -4/15 before.

OpenStudy (anonymous):

-4/12 there is no 15 in the problem sorry about that

OpenStudy (phi):

OK, so the problem *really* is \[ k - \frac{5}{12} = \frac{-4}{12} \] we have k minus 5/12 on the left side. To find k, add +5/12 to BOTH sides, like this \[ k - \frac{5}{12} + \frac{5}{12}= \frac{-4}{12} + \frac{5}{12}\] on the left side, anything minus itself is zero. that means -5/12 + 5/12 is zero. what do you get on the right side ?

OpenStudy (anonymous):

idk whats the answer if I fail this test im going to be in big trouble

OpenStudy (anonymous):

I need help plz somebody

OpenStudy (phi):

This explains how to simplify the right hand side https://www.khanacademy.org/math/arithmetic/fractions/Adding_and_subtracting_fractions/v/subtracting--fractions It takes about 2 minutes